40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Oct 26 16:20:03 2010, tarac, Notes, Laser, the laser is broken 
    Reply  Tue Nov 2 17:51:31 2010, tarac, Notes, Laser, the laser is broken 
Message ID: 378     Entry time: Tue Nov 2 17:51:31 2010     In reply to: 373
Author: tarac 
Type: Notes 
Category: Laser 
Subject: the laser is broken 

Frank opened up the laser to find any burnt mark, but found nothing and put it back, and now the laser is working.

We don't know for sure yet, what's wrong with the laser. But I'll use this opportunity to work on modification of PMC servo.

Quote:

The current 1064nm, 100 mW laser in our setup, NPRO lightwave 126 is broken. We are looking for a new one to replace it.

 

The laser stopped working when I tried to lock the cavity and saw that RCTRANSPD fluctuated a bit even after I adjust the gain setup.

So I turned the HEPA filter above the table off to see if the signal would be more stable, it was not. When I turned it back on

the laser was off. I don't think the laser and the HEPA filter are associated, but that's what happened.

 

The power output, as indicated on the laser driver is 8 mW. When I turned the laser off for 5 mins and turned it on.

The temperature of  the crystal started from ~40 C and there was power out, then, in ~10 seconds, the temperature went up to  94 C, and the power dropped to 8 mW again.

The voltage supply for TEC went up to 4V which is the maximum V for cooling.

 

I switched to the 10W laser driver, the same symptom happened again, so the problem might be the head, not the driver.

 


 

 

ELOG V3.1.3-